OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (2, 5) and (–1, 8)?

OpenStudy (anonymous):

y + 5 = 1(x + 2) y + 8 = –1(x – 1) y – 5 = –1(x – 2) y – 8 = 1(x – 2)

hero (hero):

y - y1 = m(x - x1) You're given two points. All you have to do is just insert them and solve for m: 8 - 5 = m(-1 - 2) 3 = m(-3) 3 = -3m 3m = -3 m = -3/3 m = -1 y - 5 = -1(x - 2)

OpenStudy (anonymous):

Thanks